40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 6 00:14:10 2014, Jenne, Update, LSC, ALS is iffy tonight 
    Reply  Thu Aug 7 12:34:56 2014, Jenne, Update, LSC, Suspensions not kicking? 
Message ID: 10345     Entry time: Thu Aug 7 12:34:56 2014     In reply to: 10335
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Suspensions not kicking? 

Yesterday, Q helped me look at the DACs for some of the suspensions, since Gabriele pointed out that the DACs may have trouble with zero crossings.  

First, I looked at the oplevs of all the test masses with the oplev servos off, as well as the coil drive outputs from the suspension screen which should go straight out to the DACs.  I put some biases on the suspensions in either pitch or yaw so that one or two of the coil outputs was crossing zero regularly.  I didn't see any kicks. 

Next, we turned off the inputs of the coil driver filter banks, unplugged the cable from the coil driver board to the satellite box, and put in sinusoidal excitations to each of the coils using awggui.  We then looked with a 'scope at the monitor point of the coil driver boards, but didn't see any glitches or abnormalities.  (We then put everything back to normal)

Finally, I locked and aligned the 2 arms, and just left them sitting.  The oplev servos were engaged, but I didn't ever see any big kicks. 

I am suspicious that there was something funny going on with the computers and RFM over the weekend, when we were not getting RFM connections between the vertex and the end stations, and that somehow weird signals were also getting sent to some of the optics.  Q's nuclear reboot (all the front ends simultaneously) fixed the RFM situation, and I don't know that I've seen any kicks since then, although Eric thinks that he has, at least once.  Anyhow, I think they might be gone for now.

ELOG V3.1.3-